OpenStudy (anonymous):

Given two vectors p and q, (a) write an algebraic expression for the unit vector in the direction of 3p + 2q;

OpenStudy (mathmate):

Divide any vector P by its magnitude |P| to get a unit vector.

OpenStudy (anonymous):

\[\frac{ 3p+2q }{ \sqrt{(3p+2q)^{2}} }\]

OpenStudy (anonymous):

would that =1

OpenStudy (mathmate):

Not really, p and q are vectors, so p=<a,b> q=<c,d> 3p+2q=<3a+2c, 3b+2d> and magnitude of 3p+2q |3p+2q| = |<3a+2c, 3b+2d>| =sqrt((3a+2c)^2+(3b+2d)^2)

OpenStudy (anonymous):

ok , seems good. thanks!

OpenStudy (mathmate):

In summary, if \( \textbf{P}=<a,b> \) and \( \textbf{Q}=<c,d> \), then \(3 \textbf{ P}+2\textbf{ Q} = <3a+2c, 3b+2d> \) \( \bf{ } \) and a unit vector along \(3 \textbf{ P}+2\textbf{ Q}\) is \[ \frac{<3a+2c, 3b+2d>}{\sqrt{(3a+2c)^2+(3b+2d)^2}} \]

OpenStudy (anonymous):

thanks, part two of the question is : write expression for the vector that has lenght 5 and is in opposite direction to 3p + 2q

OpenStudy (mathmate):

Multiply the previous unit vector by -5 and you'll get the required vector, since the magnitude of a unit vector is 1, and the negative sign will change it to the opposite direction.